# Large Language Monkeys00footnotetext: Title inspired by https://en.m.wikipedia.org/wiki/Infinite_monkey_theorem.: Scaling Inference Compute with Repeated Sampling

### v1

@@ Abstract
Scaling the amount of compute used to train language models has dramatically improved their capabilities. However, when it comes to inference, we often limit the amount of compute to only one attempt per problem. Here, we explore inference compute as another axis for scaling by increasing the number of generated samples. Across multiple tasks and models, we observe that coverage – the fraction of problems solved by any attempt – scales with the number of samples over four orders of magnitude. In domains like coding and formal proofs, where all answers can be automatically verified, these increases in coverage directly translate into improved performance. When we apply repeated sampling to SWE-bench Lite, the fraction of issues solved with DeepSeek-V2-Coder-Instruct increases from 15.9% with one sample to 56% with 250 samples, outperforming the single-attempt state-of-the-art of 43% which uses more capable frontier models. Moreover, using current API pricing, amplifying the cheaper DeepSeek model with five samples is more cost-effective and solves more issues than paying a premium for one sample from GPT-4o or Claude 3.5 Sonnet. Interestingly, the relationship between coverage and the number of samples is often log-linear and can be modelled with an exponentiated power law, suggesting the existence of inference-time scaling laws. Finally, we find that identifying correct samples out of many generations remains an important direction for future research in domains without automatic verifiers. When solving math word problems from GSM8K and MATH, coverage with Llama-3 models grows to over 95% with 10,000 samples. However, common methods to pick correct solutions from a sample collection, such as majority voting or reward models, plateau beyond several hundred samples and fail to fully scale with the sample budget.
@@ 1 Introduction
In addition to improving model quality, repeated sampling provides a new mechanism for minimizing LLM inference costs. When holding the total number of inference FLOPs constant, we find that on some datasets (e.g. MATH) coverage is maximized with a smaller model and more attempts, while on others (e.g CodeContests) it is better to use a larger model. We also compare API prices between DeepSeek-V2-Coder-Instruct, GPT-4o, and Claude Sonnet 3.5 in the context of solving SWE-bench Lite issues. When keeping the agent framework (Moatless Tools [62]) constant, sampling five times from the weaker and cheaper DeepSeek model solves more issues than single attempts from Claude or GPT, while also being over 3x cheaper.
@@
    We demonstrate that scaling inference compute through repeated sampling leads to large improvements in coverage across a variety tasks, models, and sample budgets. This makes it possible, and sometimes cost-effective, to amplify weaker models with many samples and outperform single attempts from more capable models. Notably, we are able to solve 56% of issues from SWE-bench Lite by sampling 250 times from DeepSeek-V2-Coder-Instruct, exceeding the single-attempt SOTA of 43%.
@@ 2 Scaling Repeated Sampling
We use the numerically stable implementation of the above formula suggested in Chen et al. [15]. Results and code will be available at https://scalyresearch.stanford.edu/pubs/large_language_monkeys/.
@@ 2.1 Repeated Sampling is Effective Across Tasks
Here, we establish that repeated sampling improves coverage across multiple tasks and a range of sample budgets. We evaluate Llama-3-8B-Instruct and Llama-3-70B-Instruct on CodeContests, MiniF2F, GSM8K, and MATH, generating 10,000 independent samples per problem. For SWE-bench Lite, we use DeepSeek-V2-Coder-Instruct [19], as the required context length of this task exceeds the limits of the Llama-3 models. As is standard when solving SWE-bench issues, we equip our LLM with a software framework that provides the model with tools for navigating through and editing codebases. In our work, we use the open-source Moatless Tools library [62]. Note that solving a SWE-bench issue involves a back-and-forth exchange between the LLM and Moatless Tools. One sample/attempt for this benchmark refers to one entire multi-turn trajectory. To minimize costs, we restrict the number of attempts per issue to 250, with all attempts made independently of one another.
@@ 2.3 Repeated Sampling Can Help Balance Performance and Cost
We also examine the dollar costs of repeated sampling when solving SWE-bench Lite issues using current API pricing. Keeping the agent framework (Moatless Tools) constant, we consider making a single attempt per issue with Claude 3.5 Sonnet and GPT-4o, as well as repeated sampling using DeepSeek-V2-Coder-Instruct. We report the average cost per issue and issue resolution rate with each approach in Table 1. While the DeepSeek model is weaker than the GPT and Claude models, it is also over 10x cheaper. In this case, repeated sampling provides a cheaper alternative to paying a premium for access to strong models while achieving a superior issue solve rate.
@@
DeepSeek-V2-Coder-Instruct 0.008 5 29.62 12 1x
GPT-4o 0.13 1 24.00 39 3.25x
Claude 3.5 Sonnet 0.17 1 26.70 51 4.25x
Table 1: Comparing API cost (in US dollars) and performance for various models on the SWE-bench Lite dataset using the Moatless Tools agent framework. When sampled more, the open-source DeepSeekCoder-V2 model can achieve the same issue solve-rate as closed-source frontier models for under a third of the price.
@@ 6 Related Work
Scaling Inference Compute: Methods that perform additional computation during inference have been successful across many areas of deep learning. Across a variety of game environments, state-of-the-art methods leverage inference-time search that examines many possible future game states before deciding on a move [12, 45, 10]. Similar tree-based methods can also be effective in combination with LLMs, allowing models to better plan and explore different approaches [58, 8, 49, 50]. Another axis for increasing LLM inference compute allows models to spend tokens deliberating on a problem before coming to a solution [57, 56, 59]. Additionally, multiple models can be ensembled together at inference time to combine their strengths [54, 14, 40, 52, 29]. Yet another approach involves using LLMs to critique and refine their own responses [39, 7].
Repeated Sampling: Previous work has demonstrated that repeated sampling can improve LLM capabilities in multiple domains. One of the most effective use cases is coding [43, 15, 34], where performance continues to scale up to a million samples and verification tools (e.g. unit tests) are often available to automatically score every candidate solution. Recently, Greenblatt [22] shows that repeated sampling is effective when solving puzzles from the ARC challenge [16], observing log-linear scaling as the number of samples increases. In chat applications, repeated sampling combined with best-of-N ranking with a reward model can outperform greedily sampling a single response [28]. In domains without automatic verification tools, existing work shows that using majority voting [55] or a trained model-based verifier [18, 38, 27] to decide on a final answer can improve performance on reasoning tasks relative to taking a single sample. Concurrent with our work, Song et al. [46] finds that using the best available sample improves LLM performance on chat, math, and code tasks, sweeping up to a max of 128 samples.
@@ A.3 MATH
We report results on $128$ randomly selected test-set problems. We sample with a temperature of $0.6$ and do not use nucleus sampling. We use the fixed $5$ few-shot example from [36] for each problem. We generate $10,000$ samples per problem. We set $512$ as the max token length for the generated solution. To grade solutions, we use the minerva_math functions from LMEval [21].
@@ A.4 GSM8K
We report results on $128$ randomly sampled test-set problems. We sample with a temperature of $0.6$ and do not use nucleus sampling. We use $5$ few-shot examples from the training set that are randomly sampled per-problem. We generate $10,000$ samples per problem. We set $512$ as the max token length for the generated solution. To grade solutions, we follow LMEval [21].
@@ B.1 Experimental Setup
For our experiments, we use DeepSeek-V2-Coder-Instruct with the Moatless Tools agent framework (at commit a1017b78e3e69e7d205b1a3faa83a7d19fce3fa6). We use Voyage AI [5] embeddings for retrieval, the default used by Moatless Tools. We make no modifications to the model or framework, using them entirely as off-the-shelf components.
@@ B.2 Test Suite Flakiness
An additional instance, astropy__astropy-6938, was flaky on some machines and not others. The authors of SWE-bench were able to reproduce the flakyness; however, we were unable to. Our prelimiary investigation indicates this specific issue is due to unpinned versions of dependencies in the docker environments that run the unit tests.
@@ Appendix D Precision Details
To calculate the Majority Vote, Reward Model + Best-of-N and Reward Model + Majority Vote metrics, we use the same $128$ problem subsets for both MATH and GSM8K datasets introduced in Section 2. Each problem corresponds to 10,000 samples for each model we test. For each verification method, we take $100$ random subsets of size $k$ and calculate the success rate using each subset. We report the mean and standard deviation across subsets in Figure 7. To calculate the Majority Vote answer, we take the plurality answer in each subset. For the Reward Model + Best-of-N, we take the answer with the highest score assigned by the reward model. For the Reward Model + Majority Vote metric, we sum the reward model score across all the samples with the same final answer, and use the final answer with the highest sum.

### v2

@@ Abstract
Scaling the amount of compute used to train language models has dramatically improved their capabilities. However, when it comes to inference, we often limit the amount of compute to only one attempt per problem. Here, we explore inference compute as another axis for scaling by increasing the number of generated samples. Across multiple tasks and models, we observe that coverage – the fraction of problems solved by any attempt – scales with the number of samples over four orders of magnitude. In domains like coding and formal proofs, where all answers can be automatically verified, these increases in coverage directly translate into improved performance. When we apply repeated sampling to SWE-bench Lite, the fraction of issues solved with DeepSeek-Coder-V2-Instruct increases from 15.9% with one sample to 56% with 250 samples, outperforming the single-attempt state-of-the-art of 43% which uses more capable frontier models. Moreover, using current API pricing, amplifying the cheaper DeepSeek model with five samples is more cost-effective and solves more issues than paying a premium for one sample from GPT-4o or Claude 3.5 Sonnet. Interestingly, the relationship between coverage and the number of samples is often log-linear and can be modelled with an exponentiated power law, suggesting the existence of inference-time scaling laws. Finally, we find that identifying correct samples out of many generations remains an important direction for future research in domains without automatic verifiers. When solving math word problems from GSM8K and MATH, coverage with Llama-3 models grows to over 95% with 10,000 samples. However, common methods to pick correct solutions from a sample collection, such as majority voting or reward models, plateau beyond several hundred samples and fail to fully scale with the sample budget.
@@ 1 Introduction
In addition to improving model quality, repeated sampling provides a new mechanism for minimizing LLM inference costs. When holding the total number of inference FLOPs constant, we find that on some datasets (e.g. MATH) coverage is maximized with a smaller model and more attempts, while on others (e.g CodeContests) it is better to use a larger model. We also compare API prices between DeepSeek-Coder-V2-Instruct, GPT-4o, and Claude Sonnet 3.5 in the context of solving SWE-bench Lite issues. When keeping the agent framework (Moatless Tools [67]) constant, sampling five times from the weaker and cheaper DeepSeek model solves more issues than single attempts from Claude or GPT, while also being over 3x cheaper.
@@
    We demonstrate that scaling inference compute through repeated sampling leads to large improvements in coverage across a variety tasks, models, and sample budgets. This makes it possible, and sometimes cost-effective, to amplify weaker models with many samples and outperform single attempts from more capable models. Notably, we are able to solve 56% of issues from SWE-bench Lite by sampling 250 times from DeepSeek-Coder-V2-Instruct, exceeding the single-attempt SOTA of 43%.
@@ 2 Scaling Repeated Sampling
We use the numerically stable implementation of the above formula suggested in Chen et al. [15]. Data and code is available at https://scalingintelligence.stanford.edu/pubs/large_language_monkeys/.
@@ 2.1 Repeated Sampling is Effective Across Tasks
Here, we establish that repeated sampling improves coverage across multiple tasks and a range of sample budgets. We evaluate Llama-3-8B-Instruct and Llama-3-70B-Instruct on CodeContests, MiniF2F, GSM8K, and MATH, generating 10,000 independent samples per problem. For SWE-bench Lite, we use DeepSeek-Coder-V2-Instruct [20], as the required context length of this task exceeds the limits of the Llama-3 models. As is standard when solving SWE-bench issues, we equip our LLM with a software framework that provides the model with tools for navigating through and editing codebases. In our work, we use the open-source Moatless Tools library [67]. Note that solving a SWE-bench issue involves a back-and-forth exchange between the LLM and Moatless Tools. One sample/attempt for this benchmark refers to one entire multi-turn trajectory. To minimize costs, we restrict the number of attempts per issue to 250, with all attempts made independently of one another.
@@ 2.3 Repeated Sampling Can Help Balance Performance and Cost
We also examine the dollar costs of repeated sampling when solving SWE-bench Lite issues using current API pricing. Keeping the agent framework (Moatless Tools) constant, we consider making a single attempt per issue with Claude 3.5 Sonnet and GPT-4o, as well as repeated sampling using DeepSeek-Coder-V2-Instruct. We report the average cost per issue and issue resolution rate with each approach in Table 1. While the DeepSeek model is weaker than the GPT and Claude models, it is also over 10x cheaper. In this case, repeated sampling provides a cheaper alternative to paying a premium for access to strong models while achieving a superior issue solve rate.
@@
DeepSeek-Coder-V2-Instruct 0.0072 5 29.62 10.8 1x
GPT-4o 0.13 1 24.00 39 3.6x
Claude 3.5 Sonnet 0.17 1 26.70 51 4.7x
Table 1: Comparing API cost (in US dollars) and performance for various models on the SWE-bench Lite dataset using the Moatless Tools agent framework. When sampled more, the open-source DeepSeek-Coder-V2-Instruct model can achieve the same issue solve-rate as closed-source frontier models for under a third of the price.
@@ 6 Related Work
Scaling Inference Compute: Methods that perform additional computation during inference have been successful across many areas of deep learning. Across a variety of game environments, state-of-the-art methods leverage inference-time search to examine many possible future game states before deciding on a move [12, 49, 10]. Similar tree-based methods can also be effective in combination with LLMs, allowing models to better plan and explore different approaches [63, 8, 53, 54]. Another axis for increasing LLM inference compute allows models to spend tokens deliberating on a problem before coming to a solution [62, 61, 64]. Additionally, multiple models can be ensembled together at inference time to combine their strengths [58, 14, 44, 56, 31]. Yet another approach involves using LLMs to critique and refine their own responses [42, 7].
Repeated Sampling: Previous work has demonstrated that repeated sampling can improve LLM capabilities in multiple domains. One of the most effective use cases is coding [47, 15, 37], where performance continues to scale up to a million samples and verification tools (e.g. unit tests) are often available to automatically score every candidate solution. Recently, Greenblatt [23] shows that repeated sampling is effective when solving puzzles from the ARC challenge [16], observing log-linear scaling as the number of samples increases. In chat applications, repeated sampling combined with best-of-N ranking with a reward model can outperform greedily sampling a single response [30]. In domains without automatic verification tools, existing work shows that using majority voting [60], prompting an LLM [19], or training a model-based verifier [18, 41, 29, 59, 35], to decide on a final answer can improve performance on reasoning tasks relative to taking a single sample. Nguyen et al. [43] finds that performing majority voting over answers that exceed a threshold length can outperform voting across all answers. Concurrent with our work, Song et al. [50] finds that using the best available sample improves LLM performance on chat, math, and code tasks, sweeping up to a max of 128 samples. Additionally, Hassid et al. [24] find that when solving coding tasks, it can be more effective to draw more samples from a smaller model than draw fewer samples from a larger one.
@@ A.3 MATH
We report results on $128$ randomly selected test-set problems. We sample with a temperature of $0.6$ and do not use nucleus sampling. We use the fixed $5$ few-shot example from [39] for each problem. We generate $10,000$ samples per problem. We set $512$ as the max token length for the generated solution. To grade solutions, we use the minerva_math functions from LMEval [22] to extract the model’s final answer. We then check correctness if the extracted answer is an exact string match to the ground truth, or if the is_equiv function from minerva_math in LMEval evaluates to true.
@@ A.4 GSM8K
We report results on $128$ randomly sampled test-set problems. We sample with a temperature of $0.6$ and do not use nucleus sampling. We use $5$ few-shot examples from the training set that are randomly sampled per-problem. We generate $10,000$ samples per problem. We set $512$ as the max token length for the generated solution. To grade solutions, we follow LMEval [22] and extract answers using a regular expression that extracts the string after the quadruple hashes. Similar to MATH, we then assess correctness by checking if the extracted answer is an exact string match to the ground truth or if is_equiv evaluates to true.
@@ B.1 Experimental Setup
For our experiments, we use DeepSeek-Coder-V2-Instruct with the Moatless Tools agent framework (at commit a1017b78e3e69e7d205b1a3faa83a7d19fce3fa6). We use Voyage AI [5] embeddings for retrieval, the default used by Moatless Tools. We make no modifications to the model or framework, using them entirely as off-the-shelf components.
@@ B.2 Test Suite Flakiness
An additional instance, astropy__astropy-6938, was flaky on some machines and not others. The authors of SWE-bench were able to reproduce the flakiness; however, we were unable to. Our preliminary investigation indicates this specific issue is due to unpinned versions of dependencies in the docker environments that run the unit tests.
@@ Appendix D Precision Details
To calculate the Majority Vote, Reward Model + Best-of-N and Reward Model + Majority Vote metrics, we use the same $128$ problem subsets for both MATH and GSM8K datasets introduced in Section 2. Each problem corresponds to 10,000 samples for each model we test. For each verification method, we take $100$ random subsets of size $k$ and calculate the success rate using each subset. We report the mean and standard deviation across subsets in Figure 7. To calculate the Majority Vote answer, we take the plurality answer in each subset (note that two answers are considered equivalent if they are exact string matches or if is_equiv evaluates to true). For the Reward Model + Best-of-N, we take the answer with the highest score assigned by the reward model. For the Reward Model + Majority Vote metric, we sum the reward model score across all the samples with the same final answer, and use the final answer with the highest sum.